OpenStudy (anonymous):

What is the value of the expression 1/3^-5 ? A. 243 B. 15 C. 1/15 D. 1/243

OpenStudy (baby_ruth):

@imqwerty

OpenStudy (anonymous):

i got d

OpenStudy (anonymous):

The value of the expression is 1/243. 3^5 stands for (3x3x3x3x3) = 243. 1/3^5 = 1/243

OpenStudy (ambi):

wrong

OpenStudy (anonymous):

well i need someone else opinion on the question then i will decide

OpenStudy (anonymous):

if im wrong its a then

OpenStudy (ambi):

\(\sf\Large\frac{1}{x^{-y}} = x^y\)

OpenStudy (anonymous):

@imqwerty

OpenStudy (ambi):

:)

imqwerty (imqwerty):

\[a^\left( -x \right)=\frac{ 1 }{ a}\]

imqwerty (imqwerty):

so \[3^\left( -5 \right)=\frac{ 1 }{ 3^5 }\]

imqwerty (imqwerty):

:)

OpenStudy (ambi):

@imqwerty the \(\large 3^{-5}\) is in the denominator though :)

OpenStudy (anonymous):

so D

OpenStudy (ambi):

It's not D :)

OpenStudy (anonymous):

wow

imqwerty (imqwerty):

\[\frac{ 1 }{ 3^\left( -5 \right) }=\frac{ 1 }{ \frac{ 1 }{ 3^5 } }=3^5=243\]

OpenStudy (ambi):

exactly :D

OpenStudy (anonymous):

oh so its A

OpenStudy (ambi):

That's what I've been saying :p

OpenStudy (anonymous):

its D though

OpenStudy (anonymous):

haha

imqwerty (imqwerty):

its A

OpenStudy (anonymous):

deffinitly A

imqwerty (imqwerty):

nd not D :)

OpenStudy (ambi):

https://www.google.com/search?q=1/3^-5 Definitely A :)

imqwerty (imqwerty):

( ͡° ͜ʖ ͡°)